Determine if the following statement is always true. If it is not, provide a counterexample. If the mathematical operation* is defined for all numbers x and y as 2x+3y, then the operation * is commutative.
Answer by tinbar(133)

x*y=2x+3y
y*x=2y+3x
clearly x*y does NOT satisfy y*x, therefore it is no commutative
let x=4, y=5
then x*y=2(4)+3(5)=8+15=23
y*x=2(5)+3(4)=10+12=22
since x*y does not equal y*x, this counterexample shows this operation definition is not commutative.